Grade 11Physical Chemistry

gaseous mixture contains 25% methane and 75% ethane by volume.
1)calculate mass of methane in mixture
2)how many total molecules are present in 10g of mixture

ritik yadav
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
mass of methane = ¼*10 = 2.5g
moles of methane = 2.5/16 = 5/32
moles of ethane = 7.5/30 = ¼
So molecules peresent will be =( ¼ + 5/32)*6.022*1023
Hope it clears.